40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 13 21:09:03 2013, manasa, Update, Green Locking, Arms ready for ALS 
    Reply  Thu Aug 15 12:30:17 2013, manasa, Update, Green Locking, Lost beat notes 
       Reply  Thu Aug 15 19:05:07 2013, manasa, Update, Green Locking, ALS out of loop noise ALS_OutLoop.pdf
Message ID: 9014     Entry time: Thu Aug 15 12:30:17 2013     In reply to: 9008     Reply to this: 9015
Author: manasa 
Type: Update 
Category: Green Locking 
Subject: Lost beat notes 

[Koji, Nic, Manasa]

Update from last night.

Koji and I realigned the green optics on the PSL to start working on the ALS.

We set on a beat note search. We couldn't find the beat note between any of the arm green transmission and the PSL green. All we could see was the beat between the X arm and the Y arm green leakage.

Since we had the beatnote between the 2 green transmission beams, we decided to scan the PSl temperature. We scanned the SLOW actuator adjust of PSL; but couldn't locate any beat note. The search will continue again today.

ELOG V3.1.3-